On the 15 March X6 Company X decides to make effective the following profit distribution corresponding to the X5 profit which was €860,000.

  • Compensation for losses from previous years of €320,000.
  • Legal reserve to comply with the Spanish legal requirement.
  • Statutory reserves of €72,000.
  • Dividend of €0.3 per share that will be made effective (paid) on the 15 April.
  • The rest will be classified as voluntary reserves.
  • The equity of Company X at that date includes: share capital of €1,200,000 (120,000 shares of €10 par value each), legal reserve of €200,000, and a voluntary reserve of €80,000.
  • We also know that on the 15 November X5 and due to the promising performance of the company during the year, a dividend was distributed and paid of €0.2 per share.
  • Dividend withholding tax is 21%.
  • Required:
  • Can you record the entry corresponding to the profit distribution on the 15 March and the entry corresponding to the dividend payment on the 15 April?
1
Expert's answer
2021-03-08T16:56:42-0500

March

Dr retained earnings 36,000

Cr dividend payable 28,440

Cr withholding tax. 7,560

April

Dr Dividend expense. 28,440

Dr withholding tax. 7,560

Cr Cash/bank 36,000

Calculation:


"Dividend payable=36,000-7560=28,440"


"Tax= .21\times36,000=7,560"

"Spain legal= 860,000\times.1=86,000"


"Dividend=.3\times120,000=36,000"